# Towards Open Federated Learning Platforms: Survey and Vision from Technical and Legal Perspectives

### 18 Jan 2024 v2

@@ 4.3.4. Batch Model Reuse Mechanisms
Amalgamation involves combining models through model parameters granularity operations, such as median (Blanchard et al., 2017; Pillutla et al., 2022) and coordinate-wise averaging with consideration of heterogeneity (McMahan et al., 2017; Li et al., 2020f), security (Sun et al., 2019), scalability (Reisizadeh et al., 2020), matching (Wang et al., 2020b; Yu et al., 2021), specificality (Gudur et al., 2020), generalizability (Qu et al., 2022a), resulting in a combination with weak separation. This reusing approach is widely used in FL works and is often referred to as ”aggregation” procedures for local models. However, in order to avoid confusion with the term ”combination,” which is frequently used interchangeably with ”aggregation” in software licenses (e.g., Artistic, GPL), we opt to use the term ”amalgamation” instead.
@@
FedAvg (McMahan et al., 2017) is the most popular model averaging method in FL with many follow-up works. For instance, Sun et al. (Sun et al., 2019) proposed applying norm thresholding of local model updates to defend against backdoor attacks. Similarly, Blanchard et al. (Blanchard et al., 2017) proposed using more robust median-based amalgamation for resilience against Byzantine behavior. Consider the ordering of parameters, Wang et al. (Wang et al., 2020b) match and average the NNs parameters layer-wise across clients, based on their similarities. Yu et al. (Yu et al., 2021) further attribute the misalignment in FL models to the non-IID training data, propose allocating an independent structure for each class and updating models through a feature paired averaging strategy. Even though amalgamation mechanisms can achieve a balance between model performance and resource efficiency by maintaining only one global model, they often rely on multiple rounds of communication to converge, which is not applicable in a query-based FL scenario. Therefore, instead of trying to directly concatenate multiple sources of models while dealing with intricate parameter mismatch, transferring the latent knowledge learned by local models to a new model is a good alternative (i.e. Federated Distillation (Jeong et al., 2018; Jin et al., 2023)).
@@
Another direction of model amalgamation is leveraging Bayesian nonparametrics to learn the shared global latent structures among local models (Yurochkin et al., 2019b, a; Lam et al., 2021). These methods, known as Model Fusion, can identify distributions of neural components across local models and only fuse the components with the same distribution, which can be regarded as a model compression between FedAvg (coordinate-wise averaging) and combination (w/o averaging). However, the model fusion strategies rely on multiple communication rounds to boost the fusion efficiency, and the model performance of one-shot fusion is even worse than that of Ensemble. Most recently, Su et al. (Su et al., 2023), inspired by null-space in continual learning (Wang et al., 2021; Kong et al., 2022), propose MA-Echo which leverages layer-wise projection matrices to preserve the original loss of local models after amalgamation. Their evaluation results present a moderate improvement in one-shot setting compared to FedAvg and ensemble. Unfortunately, this improvement is not consistently observed in multiple-round experiments. Meanwhile, to tackle the issue of catastrophic forgetting, FedPR (Feng et al., 2023) follows similar ideas to facilitate the server’s learning of visual prompts from clients for MRI reconstruction applications, but the improvement compared to FedAvg is limited even in multi-round setting.
@@
It is worth noting that our taxonomy is based on the form of the resulting model, which may not be entirely consistent with the terminology used in the technical perspective. For example, Bayes Model Averaging (BMA) (Clarke, 2003) estimates posterior probabilities of each model given the observed data, which results in a separable weighted model. Therefore, it should be classified as Combination instead of Amalgamation like FedAvg. This novel taxonomy method is useful for analyzing compatibility with licenses. For example, the coordinate-wise operations or the fusion of model parameters generate fine-grained combinations of models that are almost irreversible, which corresponds to clauses such as adapt, modify, dynamic link, etc., in software licenses.
@@
Table 4. Summary of privacy-preserving Distillation works in the field of FL. Some works are listed multiple times because they contain multiple KD procedures with different strategies. Works with naming conflicts are distinguished by subscript.
@@
FedDF (Lin et al., 2020), One-shot FL (Guha et al., 2018), FedBE (Chen and Chao, 2021), PerAda (Xie et al., 2023), FedET (Cho et al., 2022)
@@
FedFusion (Yao et al., 2019), FedKD2 (Wu et al., 2022b), MOON (Li et al., 2021b), FedNTD (Lee et al., 2022), FedMLB (Kim et al., 2022), FedCAD (He et al., 2022), FedAlign1 (Mendieta et al., 2022), FedAlign2 (Zhang et al., 2023b)
@@
FedMatch (Jeong et al., 2021), CCL (Aketi and Roy, 2024), FedProto1 (Michieli and Ozay, 2021), FedProto2 (Tan et al., 2022a)
@@
FedFusion (Yao et al., 2019), FedDistill (Jiang et al., 2020), FedKD2 (Wu et al., 2022b), MOON (Li et al., 2021b), FCCL (Huang et al., 2022b), pFedSD (Jin et al., 2022), RSCFed (Liang et al., 2022), CCL (Aketi and Roy, 2024)
@@
FedKT (Li et al., 2020b), FedMD (Li and Wang, 2019), FedAD (Gong et al., 2021), FedMD-NFDP (Sun and Lyu, 2021), FCCL (Huang et al., 2022b), FedAUX (Sattler et al., 2021), RHFL (Fang and Ye, 2022), FedKD1 (Gong et al., 2022), Cronus (Chang et al., 2021), KT-pFL (Zhang et al., 2021a), DS-FL (Itahara et al., 2021)
@@
DENSE (Zhang et al., 2022a), FedCAVE-KD (Heinbaugh et al., 2023), FedGen (Zhu et al., 2021a), FedFTG (Zhang et al., 2022d)
@@
FedKC (Wang et al., 2022), FedSSL (Fan et al., 2022)
@@
Distillation was initially proposed by Hinton et al. (Hinton et al., 2014) to transfer knowledge from a batch of independently trained neural network models (Specialists) to create a new Generalist model. Their motivation was to explore the parallelization of training of specialists and improve the efficiency of distributed NNs modeling(Dean et al., 2012). Each specialist only learns fine-grained distinctions of a subset of classes, which is very similar to the non-IID setting in FL (Li et al., 2022c). By using Knowledge Distillation (KD), we can also compress wide and deep teacher networks into lightweight student networks (Romero et al., 2015), which is promising for addressing system heterogeneity in cross-device FL (Lim et al., 2020). Therefore, it is natural to extend the KD technologies to FL filed (Jiang et al., 2020; Li et al., 2020b; Li and Wang, 2019; Wu et al., 2022b; Chen and Chao, 2021; Lin et al., 2020; Gong et al., 2021; Sun and Lyu, 2021). Actually, many recent FL works (Li and Wang, 2019; Gong et al., 2021; Sun and Lyu, 2021; Huang et al., 2022b; Fang and Ye, 2022; Gong et al., 2022; Luo et al., 2022; Sui et al., 2020; Chang et al., 2021; He et al., 2020a; Zhang et al., 2021a; Itahara et al., 2021; Aketi and Roy, 2024; Michieli and Ozay, 2021; Tan et al., 2022a) solely leverage KD without following the model averaging paradigm of FedAvg, we leave this discussion for later.
@@
Despite directly retraining a Generalist model through knowledge distillation (KD), an alternative approach is to construct an ensemble of knowledge. For example, Furlanello et al. (Furlanello et al., 2018) consecutively generate student models with the guidance of knowledge distilled from earlier generations and find that the ensemble of multiple generations of internal models achieves state-of-the-art performance. Dvornik et al. (Dvornik et al., 2019) leverage the distilled knowledge from each learner to encourage cooperation and prediction diversity within the population, which leads to better ensemble results. In the context of FL, the main advantage of distillation is the decoupling between knowledge distillation and knowledge learning. This allows us to split the model architecture for the purpose of system heterogeneity and efficiency (Vepakomma et al., 2019; Thapa et al., 2022). Moreover, the well-learned knowledge from clients only needs to be communicated once (Gong et al., 2021, 2022), while the server can perform multiple epochs of local training to complete the transfer. The drawback of KD is that it is data-dependent and the shared knowledge may be extracted from local sensitive data, which exposes a new attack surface for potential model inversion attacks (Kim et al., 2020; Fredrikson et al., 2015). To mitigate this issue, some efforts (Wang et al., 2022; Fan et al., 2022) have been made to add differential privacy noise (Dwork, 2006) to the shared content.
@@
In general, there are three mechanisms for avoiding the sharing of sensitive knowledge. First, push the KD procedure to the server-side, where the knowledge of local models is transferred to the global model through a validation set (Sui et al., 2020) or unlabeled dataset (Lin et al., 2020; Guha et al., 2018; Chen and Chao, 2021; Xie et al., 2023; Cho et al., 2022) held by the server. Second, we can keep the KD procedure at the client-side, allowing the knowledge of the global model (Yao et al., 2019; Wu et al., 2022b; Li et al., 2021b; Lee et al., 2022; Mendieta et al., 2022; Kim et al., 2022; He et al., 2022; Zhang et al., 2023b), other clients’ models (Jeong et al., 2021; Aketi and Roy, 2024; Michieli and Ozay, 2021; Tan et al., 2022a) or self-model (Yao et al., 2019; Jiang et al., 2020; Wu et al., 2022b; Li et al., 2021b; Huang et al., 2022b; Jin et al., 2022; Liang et al., 2022) to be transferred based on the local training data. In the above two strategies, only model parameters are exchanged in the training network, which means they can provide the same level of privacy protection as traditional FL.
@@
The last mechanism is to assume that a public unlabeled dataset, which does not contain sensitive information, can be accessed by both the server and clients for KD (Li et al., 2020b; Li and Wang, 2019; Gong et al., 2021; Sun and Lyu, 2021; Huang et al., 2022b; Sattler et al., 2021; Fang and Ye, 2022; Gong et al., 2022; Chang et al., 2021; Zhang et al., 2021a; Itahara et al., 2021). Sharing these extracted contents will not raise any privacy concerns, and only minimal communication is generated during KD for the purpose of aligning sample IDs. In cases where such public datasets are not available on the server, a recent approach called Data-Free Knowledge Distillation (DFKD) has been proposed by Lopes et al. (Lopes et al., 2017). This approach regenerates batches of data based on layer activation statistics or spectrum coefficients collected during training phase, and this synthetic data is then used for distillation. DENSE (Zhang et al., 2022a) is the first attempt to extend DFKD to FL. It leverages the ensemble of local models to guide the training of a data generator on the server, and the generated data is then used to distill the knowledge from local models to the global model. FedCAVE-KD (Heinbaugh et al., 2023) leverages locally trained conditional autoencoders (CVAEs) (Kingma and Welling, 2014) to generate samples based on the data distribution of clients. These CVAEs are sent to server used to construct a global generator via KD, which will later provide synthetic training data for the global discriminator.
@@
It is worth noting that in the query-based FL setting, direct access to the original data is not available, thus the second mechanism mentioned earlier cannot be directly applied. A circumvention method is to train a generator following the inspiration of DFKD. Fortunately, this is praticable if the workflow and history information of modeling are tracked and queryable, as we advocated in Section 4.2.2. Recalling that, as shown in Fig. 4(b), Generation is the last category in our taxonomy. Actually, such a hybrid model reuse strategy is quite common in FL. For example, the previously mentioned DENSE (Zhang et al., 2022a) incorporates three model reuse mechanisms: Combination (creating an ensemble), Generation (generating synthetic data), and Distillation. Therefore, our taxonomy can cover traditional FL works, such as FedAvg and MOON (Li et al., 2021b), as well as the broad sense FL, including Federated Distillation (Jeong et al., 2018; Jin et al., 2023) and Ensemble Learning (Shi et al., 2023; Wang et al., 2023b). We provide a comparison of these hybrid works in Section 4.3.5. The summarize of above privacy-perserving KD works is given in Table. 4.
@@
Table 5. Classification of Generation works in the field of FL. Some of the works are also listed in Table 4 because they utilize hybrid model reuse mechanisms.
@@
FedSage+ (Zhang et al., 2021c), GFL (Cheng et al., 2023),
FRD (Cha et al., 2019)
Fed-ZDA (Hao et al., 2021), FOSTER (Yu et al., 2023), FedDG (Liu et al., 2021a), DynaFed (Pi et al., 2022), SDA-FL (Li et al., 2022e), FedMix (Yoon et al., 2021),
FedCAVE-Ens (Heinbaugh et al., 2023)
FedDISC (Yang et al., 2023), SemiFL (Diao et al., 2022)
@@
DENSE (Zhang et al., 2022a), FedBE (Chen and Chao, 2021),
FedDyn (Jin et al., 2023), FedZKT (Zhang et al., 2022e)
FedGen (Zhu et al., 2021a), FedFTG (Zhang et al., 2022d), FD+FAug (Jeong et al., 2018),
FedCAVE-KD (Heinbaugh et al., 2023)
FedSSL (Fan et al., 2022)
Generation is designed to generate synthetic samples that resemble the original data distribution by building a probabilistic model (Geyer, 1992) or deep learning model (Kingma and Welling, 2014; Goodfellow et al., 2020; Cao et al., 2022) that can capture the underlying distribution pattern and latent structure of original data. Generally speaking, generation techniques can be classified into three categories: data-level, probabilistic, and representation-based approaches. Data-level approaches involve sample granularity operations such as interpolation (Chawla et al., 2002; Zhang et al., 2018) and augmentation (Wong et al., 2016) to generate synthetic features based on the original feature space of data and share. Even though these methods are training-free and easy to implement, they cannot be directly applied to the FL setting due to privacy concerns. Recent proposed FedMix (Yoon et al., 2021) aims to alleviate the negative effect of non-IID data by using mixup (Zhang et al., 2018), where the average of local data is linearly interpolated with the training data to generate augmented samples. However, the potential risk of data leakage when sharing the mixup data is not comprehensively evaluated in the original work.
Probabilistic approachs aim to estimate the real data distribution using probabilistic method. For example, Markov Chain Monte Carlo (MCMC) (Geyer, 1992) methods construct a Markov Chain that converges to the desired target distribution by iteratively proposing new states based on the current state of the chain and acceptance probabilities. Gaussian Mixture Models (GMM) assumes the data is generated from a mixture of Gaussian distributions and can generate new samples by sampling from the learned distributions. To generate the high-dimensional structured data, representation-based approaches try to reconstruct the data from laten feature space. For instance, Variational Autoencoders (VAEs) (Kingma and Welling, 2014) learn the distribution of the latent representation space given the observed data and then use a decoder network to reconstruct data based on sampled latent representations. Generative Adversarial Networks (GANs) (Goodfellow et al., 2020) train a generator network to produce samples that resemble realistic data by optimizing an adversarial objective against a discriminator network.
Compare to the other model reuse mechanisms, generation has three unique advantages. The first advantage is visualization and verification. Unlike the extracted knowledge in distillation, the quality of generated content can be visualized and validated by humans. This capability aids in assessing the contributions made by participants in terms of generating valuable content. Second, the flexibility of generation methods allow us to generate data with any desired amount or class, which enables more effective handling of imbalanced (Chawla et al., 2002) and non-IID (Zhang et al., 2022d) data. The third advantage is multi-format sharing. Participants have the freedom to choose the form of their contributions. For example, they can upload the learned generative models in source code (e.g., Stable Diffusion (Rombach et al., 2022), GPT-2 (Radford et al., 2019)) or binary form, upload synthetic data, or provide model inference APIs like ChatGPT. This sharing policy can greatly empower the model community in open FL platforms, fostering collaboration and knowledge sharing.
Given the aforementioned advantages, generation methods have been extensively studied in the field of FL. As summaried in Table. 5, these works can be classified into two main categories: generation for training (Zhang et al., 2021c; Cheng et al., 2023; Hao et al., 2021; Cha et al., 2019; Yu et al., 2023; Heinbaugh et al., 2023; Yang et al., 2023; Liu et al., 2021a; Pi et al., 2022; Li et al., 2022e; Diao et al., 2022), generation for KD (Zhang et al., 2022a; Chen and Chao, 2021; Zhu et al., 2021a; Zhang et al., 2022d; Jeong et al., 2018; Jin et al., 2023; Heinbaugh et al., 2023; Zhang et al., 2022e; Fan et al., 2022), and serving three purposes: enriching the training set (Zhang et al., 2022a; Chen and Chao, 2021; Zhang et al., 2021c; Cheng et al., 2023; Cha et al., 2019; Jin et al., 2023; Zhang et al., 2022e), improving generalization ability (Zhu et al., 2021a; Zhang et al., 2022d; Hao et al., 2021; Jeong et al., 2018; Yu et al., 2023; Heinbaugh et al., 2023; Liu et al., 2021a; Pi et al., 2022; Li et al., 2022e), and enabling semi-supervised learning (Yang et al., 2023; Diao et al., 2022; Fan et al., 2022). As an example of generation for training, FedSage+ (Zhang et al., 2021c) trains a missing neighbors generator to mend the links between cross-subgraph nodes, thereby increasing the connectivity of local data and benefiting from this collaboration across clients. Previous mentioned FedCAVE-KD (Heinbaugh et al., 2023) is an example of generation for KD, where locally trained CVAEs and local label distributions are uploaded to the server for DFKD, ensuring privacy while also enhancing generalization of global model. Another example is FedGen (Zhu et al., 2021a), where the generator is maintained by the server and sent to clients in each round. The generator has knowledge about the global view of the data distribution, which is used to KD into the local models, thereby enhancing their generalizability. The last application of generation is in semi-supervised learning, which is a common real-world scenario where the client data is unlabeled. For example, FedDISC (Yang et al., 2023) leverages the average and cluster centroids of hidden representations across pseudo-labels as input to a pre-trained diffusion model, aiming to generate high-quality samples for training.
In fact, the three purposes of generation correspond to three types of data heterogeneity in FL (Li et al., 2022c): Quantity Skew, Label Distribution Skew (non-IID), and Missing Labels, which are challenging to address with traditional model amalgamation methods. The hybrid model reusing strategies, which leverage each other’s strengths, have become a common paradigm in recent FL studies (Cheng et al., 2023; Xie et al., 2023; Yu et al., 2023; Jin et al., 2023; Heinbaugh et al., 2023; Yang et al., 2023; Zhang et al., 2023b). Therefore, in the next section, we will summarize the popular hybrid model reusing studies in FL and then filter the studies that are suitable for query-based FL platforms.

### 29 Feb 2024 v3

@@ V-B Amalgamation
Amalgamation involves combining models through model parameters granularity operations, such as median [122, 123] and coordinate-wise averaging with consideration of heterogeneity [8, 10], security [124], scalability [125], matching [126, 127], specificality [128], generalizability [129], resulting in a combination with weak separation. This reusing approach is widely used in FL works and is often referred to as "aggregation" procedures for local models. However, in order to avoid confusion with the term "combination," which is frequently used interchangeably with "aggregation" in software licenses (e.g., Artistic, GPL), we opt to use the term "amalgamation" instead.
@@
FedAvg [8] is the vanilla model averaging method in FL with many follow-up works. For instance, Sun et al. [124] introduced applying norm thresholding of local model updates to defend against backdoor attacks. Similarly, Blanchard et al. [122] proposed using more robust median-based amalgamation for resilience against Byzantine behavior. Consider the ordering of parameters, Wang et al. [126] match and average the NNs parameters layer-wise across clients, based on their similarities. Yu et al. [127] further attribute the misalignment in FL models to the non-IID training data, and propose allocating an independent structure for each class and updating models through a feature paired averaging strategy.
While amalgamation mechanisms can achieve a balance between model performance and resource efficiency by maintaining only one global model, they often rely on multiple rounds of communication to converge, which is not applicable in a query-based FL scenario. Therefore, instead of trying to directly concatenate multiple sources of models while dealing with intricate parameter mismatch, transferring the latent knowledge learned by local models to a new model is a good alternative (i.e. Federated Distillation [130, 131]).
@@
Another direction of model amalgamation is leveraging Bayesian nonparametrics to learn the shared global latent structures among local models [132, 133, 134]. These methods, known as Model Fusion, can identify distributions of neural components across local models and only fuse the components with the same distribution, which can be regarded as a model compression between FedAvg (coordinate-wise averaging) and combination (w/o averaging). However, the model fusion strategies rely on multiple communication rounds to boost the fusion efficiency, and the model performance of one-shot fusion is even worse than that of Ensemble. Most recently, Su et al. [135], inspired by null-space in continual learning [136, 137], propose MA-Echo which leverages layer-wise projection matrices to preserve the original loss of local models after amalgamation. Their results present a moderate improvement in one-shot setting compared to FedAvg and ensemble.
Unfortunately, this improvement is not consistently observed in multiple-round experiments. Meanwhile, to tackle the issue of catastrophic forgetting, FedPR [138] follows similar ideas to facilitate the server’s learning of visual prompts from clients for MRI reconstruction applications, but the improvement is limited even in multi-round setting.
@@
It is worth noting that our taxonomy is based on the form of the resulting model, which may not be entirely consistent with the terminology used in the technical perspective. For example, Bayes Model Averaging (BMA) [114] estimates posterior probabilities of each model given the observed data, which results in a separable weighted model. Therefore, it should be classified as Combination instead of Amalgamation like FedAvg. This novel taxonomy method is useful for analyzing compatibility with licenses. For example, the coordinate-wise operations or the fusion of model parameters generate fine-grained combinations of models that are almost irreversible, which corresponds to clauses such as adapt, modify, dynamic link, and etc., in software licenses.
@@
TABLE III: Summary of privacy-preserving Distillation works in the field of FL. Some works are listed multiple times because they contain multiple KD procedures with different strategies. Works with naming conflicts are distinguished by subscript.
@@
FedDF [140], One-shot FL [141], FedBE [142], PerAda [143], FedET [144]
@@
FedFusion [145], FedKD2 [146], MOON [12], FedNTD [147], FedMLB [148], FedCAD [149], FedAlign1 [150], FedAlign2 [151]
@@
FedMatch [152], CCL [153], FedProto1 [154], FedProto2 [155]
@@
FedFusion [145], FedDistill [156], FedKD2 [146], MOON [12], FCCL [157], pFedSD [158], RSCFed [159], CCL [153]
@@
FedKT [160], FedMD [161], FedAD [162], FedMD-NFDP [163], FCCL [157], FedAUX [164], RHFL [165], FedKD1 [166], Cronus [167], KT-pFL [168], DS-FL [169]
@@
DENSE [170], FedCAVE-KD [171], FedGen [172], FedFTG [173]
@@
FedKC [174], FedSSL [175]
V-C Distillation
@@
Distillation was initially proposed by Hinton et al. [77] to transfer knowledge from a batch of independently trained neural network models (Specialists) to create a new Generalist model. Their motivation was to explore the parallelization of training of specialists and improve the efficiency of distributed NNs modeling[176]. Each specialist only learns fine-grained distinctions of a subset of classes, which is very similar to the non-IID setting in FL [177]. By using Knowledge Distillation (KD), we can also compress wide and deep teacher networks into lightweight student networks [178], which is promising for addressing system heterogeneity in cross-device FL [179]. Therefore, it is natural to extend the KD technologies to FL filed [146]. Many recent FL works [153, 180, 155] solely leverage KD without following the model averaging paradigm of FedAvg, we leave this discussion for later.
@@
Despite directly retraining a Generalist model through KD, an alternative approach is to construct an ensemble of knowledge. For example, Furlanello et al. [181] consecutively generate student models with the guidance of knowledge distilled from earlier generations and find that the ensemble of multiple generations of internal models achieves state-of-the-art performance. Dvornik et al. [182] leverage the distilled knowledge from each learner to encourage cooperation and prediction diversity within the population, which leads to better ensemble results. In the context of FL, the main advantage of distillation is the decoupling between KD and knowledge learning. This allows us to split the model architecture for the purpose of system heterogeneity and efficiency [183, 184]. Moreover, the well-learned knowledge from clients only needs to be communicated once [162, 166], while the server can perform multiple epochs of local training to complete the transfer.
The drawback of KD is that it is data-dependent and the shared knowledge may be extracted from local sensitive data, which exposes a new attack surface for potential model inversion attacks [185, 186]. To mitigate this issue, some efforts [174, 175] have been made to add differential privacy noise [45] to the shared content.
@@
In general, there are three mechanisms for avoiding the sharing of sensitive knowledge. First, push the KD procedure to the server-side, where the knowledge of local models is transferred to the global model through a validation set [139] or unlabeled dataset [140, 141, 142, 143, 144] held by the server. Second, we can keep the KD procedure at the client-side, allowing the knowledge of the global model [145, 146, 12, 147, 150, 148, 149, 151], other clients’ models [152, 153, 154, 155] or self-model [145, 156, 146, 12, 157, 158, 159] to be transferred based on the local training data. In the above two strategies, only model parameters are exchanged in the training network, which means they can provide the same level of privacy protection as traditional FL.
@@
The last mechanism is to assume that a public unlabeled dataset, which is non-sensitive, is accessible by both the server and clients for KD [160, 161, 162, 163, 157, 164, 165, 166, 167, 168, 169]. Sharing these extracted contents will not raise any privacy concerns, and only minimal communication is generated during KD for the purpose of aligning sample IDs. In cases where such public datasets are not available on the server, a recent approach known as Data-Free Knowledge Distillation (DFKD) [187] regenerates batches of data based on layer activation statistics or spectrum coefficients collected during training phase. Then, this synthetic data is used for distillation. DENSE [170] is the first attempt to extend DFKD to FL. It leverages the ensemble of local models to guide the training of a data generator on the server, and the generated data is then used to distill the knowledge from local models to the global model. FedCAVE-KD [171] leverages locally trained conditional autoencoders (CVAEs) [188] to generate samples based on the data distribution of clients. These CVAEs are sent to server used to construct a global generator via KD, which will later provide synthetic training data for the global discriminator.
@@
It is worth noting that in the query-based FL setting, direct access to the original data is not available, thus the second mechanism mentioned earlier cannot be directly applied. A circumvention method is to train a generator following the inspiration of DFKD. Fortunately, this is praticable if the workflow and history information of modeling are tracked and queryable, as we advocated in §IV-B. Recalling that, as shown in Fig. 5(b), Generation is the last category in our taxonomy. Actually, such a hybrid model reuse strategy is quite common in FL. For example, the previously mentioned DENSE [170] incorporates three model reuse mechanisms: Combination (creating an ensemble), Generation (generating synthetic data), and Distillation. Therefore, our taxonomy can cover traditional FL works, such as FedAvg and MOON [12], as well as the broad sense FL, including Federated Distillation [130, 131] and Ensemble Learning [189, 190]. We provide a comparison of these hybrid works in §V-E. The summarization of above privacy-perserving KD works is given in TABLE III.
V-D Generation
Generation is designed to generate synthetic samples that resemble the original data distribution by building a probabilistic model [191] or deep learning model [188, 192, 84] that can capture the underlying distribution pattern and latent structure of original data. Generally speaking, generation techniques can be classified into three categories: data-level, probabilistic, and representation-based approaches. Data-level approaches involve sample granularity operations such as interpolation [193, 194] and augmentation [195] to generate synthetic features based on the original feature space of data and share. Even though these methods are training-free and easy to implement, they cannot be directly applied to the FL setting due to privacy concerns. Recent proposed FedMix [196] aims to alleviate the negative effect of non-IID data by using mixup [194], where the average of local data is linearly interpolated with the training data to generate augmented samples. However, the potential risk of data leakage when sharing the mixup data is not comprehensively evaluated in the original work.
Probabilistic approachs aim to estimate the real data distribution using probabilistic method. For example, Markov Chain Monte Carlo (MCMC) [191] methods construct a Markov Chain that converges to the desired target distribution by iteratively proposing new states based on the current state of the chain and acceptance probabilities. Gaussian Mixture Models (GMM) assumes the data is generated from a mixture of Gaussian distributions and can generate new samples by sampling from the learned distributions. To generate the high-dimensional structured data, representation-based approaches try to reconstruct the data from laten feature space. For instance, Variational Autoencoders (VAEs) [188] learn the distribution of the latent representation space given the observed data and then use a decoder network to reconstruct data based on sampled latent representations. Generative Adversarial Networks (GANs) [192] train a generator network to produce samples that resemble realistic data by optimizing an adversarial objective against a discriminator network.
Compared to the other model reuse mechanisms, generation has three unique advantages. The first advantage is visualization and verification. Unlike the extracted knowledge in distillation, the quality of generated content can be visualized and validated by humans. This capability aids in assessing the contributions made by participants in terms of generating valuable content. Second, the flexibility of generation methods allow us to generate data with any desired amount or class, which enables more effective handling of imbalanced [193] and non-IID [173] data. The third advantage is multi-format sharing. Participants have the freedom to choose the form of their contributions. For example, they can upload the learned generative models in source code (e.g., Stable Diffusion [91], GPT-2 [78]) or binary form, upload synthetic data, or provide model inference APIs like ChatGPT. This sharing policy can greatly empower the model community in open FL platforms, fostering collaboration and knowledge sharing.
Given the aforementioned advantages, generation methods have been extensively studied in the field of FL. As summaried in TABLE IV, these works can be classified into two main categories: generation for training [197, 198, 199, 200, 201, 171, 202, 203, 204, 205, 206], generation for KD [170, 142, 172, 173, 130, 131, 171, 207, 175], and serving three purposes: enriching the training set [170, 142, 197, 198, 200, 131, 207], improving generalization ability [172, 173, 199, 130, 201, 171, 203, 204, 205], and enabling semi-supervised learning [202, 206, 175]. As an example of generation for training, FedSage+ [197] trains a missing neighbors generator to mend the links between cross-subgraph nodes, thereby increasing the connectivity of local data and benefiting from this collaboration across clients. Previously mentioned FedCAVE-KD [171] is an example of generation for KD, where locally trained CVAEs and local label distributions are uploaded to the server for DFKD, ensuring privacy while also enhancing generalization of global model. Another example is FedGen [172], where the generator is maintained by the server and sent to clients in each round. The generator has knowledge about the global view of the data distribution, which is used to KD into the local models, thereby enhancing their generalizability. The last application of generation is in semi-supervised learning, which is a common real-world scenario where the client data is unlabeled. For example, FedDISC [202] leverages the average and cluster centroids of hidden representations across pseudo-labels as input to a pre-trained diffusion model, aiming to generate high-quality samples for training.
In fact, the three purposes of generation correspond to three types of data heterogeneity in FL [177]: Quantity Skew, Label Distribution Skew (non-IID), and Missing Labels, which are challenging to address with traditional model amalgamation methods. The hybrid model reusing strategies, which leverage each other’s strengths, have become a common paradigm in recent FL studies [198, 143, 201, 131, 171, 202, 151]. Therefore, in the next section, we will summarize the popular hybrid model reusing studies in FL and then filter the studies that are suitable for query-based FL platforms.
@@ V-D Generation
TABLE IV: Classification of Generation works in the field of FL. Some of the works are also listed in TABLE III because they utilize hybrid model reuse mechanisms.
@@
FedSage+ [197], GFL [198],
FRD [200]
Fed-ZDA [199], FedDG [203], FOSTER [201], DynaFed [204], SDA-FL [205], FedMix [196],
FedCAVE-Ens [171]
FedDISC [202],
SemiFL [206]
@@
DENSE [170], FedBE [142],
FedDyn [131], FedZKT [207]
FedGen [172], FedFTG [173], FD+FAug [130], FedCAVE-KD [171]
FedSSL [175]
V-E Hybrid Model Reusing and Model Licenses
Following the taxonomy we introduced in §V, it can be observed that almost all FL studies can be regarded as a permutation of four model reuse mechanisms: Combination, Amalgamation, Distillation, and Generation. To enhance our understanding of current model reusing studies and identify methods applicable for constructing an open FL platform, we provide a comprehensive summary in TABLE V. We employ different colors and fonts in TABLE V to emphasize the distinctions among studies, while certain processes have been omitted without ambiguity. In addition, we have listed the main goals of each study, and only those with explicit designs, experiments, or proof are counted. Please refer to the table caption for the explanation of our denotations. As an example, the process of RSCFed [159] involves the following steps: ① KD, the knowledge is extracted as the softmax values from the self model based on local private data (ref. TABLE III) and performed at the client-side; ② Model exponential moving averaging performed at the client-side; ③ Simple Model Averaging performed two times at the server-side. These processes are repeated across multiple communication rounds until completion. In this way, we can categorize these works and make intuitive comparisons between them‡IV-A.
Through disassembling hybrid model reusing into the four mechanisms, we can further analyze the corresponding clauses in model licenses perspective for FL studies. Then we can easily identify the applicable clauses for different reusing mechanisms‡IV-B and analyze potential licenses conflicts [^3] We developed ModelGo to implement this idea..